LHC 2012 - latest news

Week 48: Machine coordinators: Gianluigi Arduini, Eva Barbara Holzer

MD Coordination: Frank Zimmermann, Giulia Papotti

Goal of this week: MD & Luminosity production.

Date

Summary

Link to report from morning meeting

Mon 3 December   08:30 meeting
MD summary
Sun 2 December
  • 1:23 beam dump in adjust: Slow losses close to roman pots (MQML 6R5) during collapse of IP1/5 separation bumps
  • 3:55 16:05 stable beams fill 3360 (OP dump); Initial luminosity up to 7.2E33 cm-2s-1; ALICE at ~0.9E31 cm-2s-1. Integrated lumi: 172 pb -1 in 12 hours
  • 8:52 beam dump in Adjust: UFO like loss (~0.6ms long) at BLMQI.06L5.B2E21_MQML_XRP
  • 20:34 beam dump during Ramp; Loss B2 (possibly a long UFO or a slow loss, rising for 3ms) starting in cell 4L6 triggering BLM on TCDQM.4L6
  • Satellite enhancement reversed again and re-filling
  • 23:27 stable beam fill 3363; Initial luminosity up to 7.1E33 cm-2s-1
09:00 meeting
Sat 1 December
  • 01:45 Loss of communication with RF FGC. Abort gap populaton increasing.
  • 02:57 Access required to solve problem RF problem. OP dump after abort gap cleaning. End of fill #3350. 120 pb-1 in 7.7 h.
  • 04:49 Access finished (RF control crate power supply changed). Problem to move the TED in TI2.
  • 08:12 Stable beams #3351. Initial luminosity up to 6.9E33 cm-2s-1. ALICE at ~1E31.
  • 16:00 OP dump. End of fill #3351. 120 pb -1 in 7.7 hours.
  • Attempt to put in place compensatory measures for the unavailability of the TI2 TED proposed by the DSO. Not possible.
  • New extraction bump configuration in place in the PS (extraction conditions at the extraction septum in the PS are constant)
  • 19:30 injecting and steering transfer lines
  • 20:30-22:30 Losses at injection and LHCb trigger due satellites on the injection kicker pulse edge. Discovered problem in the copy of the RF and kicker synchronization settings. Go back to standard filling scheme.
  • 23:30 beam dump during injection by RF crowbar event on M1B1 --> reset
09:00 meeting
Fri 30 November
  • 00:10 Access finished
  • 02:16 STABLE BEAMS #3347. Initial luminosity ~7.3e33 cm-2s-1. Initial ALICE luminosity still too low (5.5e30cm-2s-1)
  • 12:07 OP dump. End of fill #3347. ~156 pb-1 in 9.8 h
  • 14:00-17:00 Attempt of injecting high brightness beam failed due to timing issue.
  • 17:00 Filling for physics #3350
  • 19:10 STABLE BEAMS #3350.Initial luminosity 6.7E33 cm-2s-1.
8:30 meeting
Thu 29 November
  • 04:00 End of MD 4
  • 04:0007:30 Test ramp and asynch. dump test in collision
  • 07:30 - 10:30 IR8 aperture measurement
  • 10:30-15:20 Access.
  • 15:00 - 16:30. 50 ns beam checked by RF expert in PS and satellite enhancement ON
  • 16:30 ready to inject. No beam from SPS (coming out of MD)
  • 17:40 Finally beam available. Steering of the injection lines.
  • 18:40-21:45 Operational development: HOM heating vs. bunch length
  • 22:48 Trip of RCBXH2.L5 while filling for physics. Access required for power converter repair.
MD summary
8:30 meeting
Wed 28 November
  • MD4
MD summary
08:30 meeting
Tue 27 November
  • MD4
MD summary
Mon 26 November
  • fill 3323 dump at 06:00 Monday morning for start of MD program
  • IR8 aperture MD postponed due to QPS problem requiring access
08:30 meeting